OpenStudy (anonymous):

which of the binomials below is a factor of this trinomial? : x^2+3x-28 . . . A. x-7 B. x^2+7 C. x+4 D. x+7

OpenStudy (anonymous):

What do u think?

OpenStudy (anonymous):

i dont know.. what is it?

OpenStudy (anonymous):

i give you the hint

OpenStudy (anonymous):

okay what

OpenStudy (anonymous):

X+4?

OpenStudy (anonymous):

x^2+3x-28 = x^2 +7x-4x -28 = x(x+7)-4(x+7) can you do it now

OpenStudy (anonymous):

no

OpenStudy (anonymous):

X+7?

OpenStudy (anonymous):

yeeees

OpenStudy (anonymous):

ohhhhhh...thanks.

OpenStudy (anonymous):

welcomeeeee

OpenStudy (anonymous):

okay soo what about this one... haha. x^2+4x-60. A. x+6. B. x-10. C. x-6. D. x^2-6

OpenStudy (anonymous):

ok x^2 +10x-6x-60 x(x+10)-6(x+10) now guesssssssss

OpenStudy (anonymous):

okay.. um B? x-10.. lol

OpenStudy (anonymous):

no

OpenStudy (anonymous):

x+6?!?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

okay what i really dont knomww.. hah

OpenStudy (anonymous):

i do one step more (x-6)(x+10) guess it now

OpenStudy (anonymous):

x-6?

OpenStudy (anonymous):

yeeeeeeeeeeeeeeeeeeeeeee

OpenStudy (anonymous):

x-6 is the right answer

OpenStudy (anonymous):

okaay yay! so whats this one.. x^2+2x-35... A. x+5 B. x-7 C. x^2+7. D. x-5.

OpenStudy (anonymous):

x^2+2x-35 = x^2 +7x -5x -35 = x(x+7)-5(x+7)-----> do it now

OpenStudy (anonymous):

x-5? or x-7?

OpenStudy (anonymous):

x-5

OpenStudy (anonymous):

how you guess x-7 its wrong

OpenStudy (anonymous):

x^2+19x+84.... A. x^2+12. B. x-12 C. x+12. D. x+84

OpenStudy (anonymous):

this time you have to guess from only one step hint x^2+19x+84 = x^2 +12x+7x+84 do it now

OpenStudy (anonymous):

@blondechic4